OpenStudy (anonymous):

4. The value of y is inversely proportional to the value of x. When y=42, x=6. What is the value of y when x=9 ? A. 252 B. 63 C. 28 D. 45

OpenStudy (anonymous):

C 28 y=k/x (inverse) 42=k/6 k=42*6 y=(42*6)/9 y=28